Winebulance Co. purchased a new van on 1-1-2019 for Estimated salvage value
Estimated useful life
Estimated total miles over the useful life
Actual miles driven in 2019 Actual miles driven in 2020
84 000 4 000
4 years 160 000 miles 37 000 miles 45 000 miles
INSTRUCTIONS: Compute depreciation expense for the years 2019 and 2020 using each of the following methods:
1. Straight line
2. Units of Activity
3. Double declining balance
Answer #1

Answer-1)- The depreciation expense for 2019= $20000.

The depreciation expense for 2020= $20000.

Explanation- Straight line Method-

= Cost of asset- Salvage value of asset/No. of useful life (years)

=($84000-$4000)/4 years

=$80000/ 4 years

= $20000         

Year 2019 depreciation expense = $20000

Year 2020 depreciation expense = $20000

2)- The depreciation expense for 2019= $18500.

The depreciation expense for 2020= $22500.

Explanation- Units of activity method:-Annual depreciation expense per unit-

=Cost – salvage /Total miles driven

=($84000-$4000)/160000 miles

=$0.50 per mile

Depreciation expense in year 2019= Depreciation expense per hour* Hours used

                                                =$0.50 per mile *37000 miles

                                                 =$18500

Depreciation expense in year 2020= Depreciation expense per hour* Hours used

                                                =$0.50 per mile *45000 miles

                                                 =$22500

3)- The depreciation expense for 2019 = $33600.

The depreciation expense for 2020 = $25200.

Explanation- Double Declining balance depreciation is calculated using the following formula:

Depreciation = Depreciation Rate * Book Value of Asset

Depreciation rate is given by the following formula:

Depreciation Rate = Accelerator *Straight Line Rate

Straight-line Depreciation Rate = 1/4 = 0.25 = 25%
Declining Balance Rate = 2*25% = 50%

Diminishing balance method – Cost of equipment*Depreciation rate

Depreciation expense Year 2019 = $84000*40%

= $33600

Book value at the end of year 2019 = $84000-$33600 = $50400

Year 2019 Depreciation expense = $50400*50%= $25200